Anton is twice as old as his wife Marie. Eight years ago, Anton was eight years younger than three time's Marie's age. How old is Marie now?
Answer by stanbon(75887)

Anton is twice as old as his wife Marie. Eight years ago, Anton was eight years younger than three time's Marie's age. How old is Marie now?
-------------------
Equations:
a = 2m
(a-8) = 3(m-8)-8
----
Rearrange:
a = 2m
a-8 = 3m-32
---
Substitute for "a" and solve for "m":
2m-8 = 3m-32
m = 24 (Marie's age now)
----
Solve for "a":
a = 2m
a = 48 (Anton's age now)
===========
Cheers,
Stan H.
=============